OpenStudy (anonymous):

integrate (X^2-x +1) sin x dx

OpenStudy (anonymous):

\[\int\limits( x^2 +x+1) \sin x dx\]

OpenStudy (anonymous):

Looks hella like you could use integration by parts.

OpenStudy (anonymous):

if i expand this, I get \[\int\limits (x^2 \sin x + x \sin x + \sin x) dx\] then I sould just do \[\int\limits u dv= uv- \int\limits v du \] for each

OpenStudy (anonymous):

That is a way to do it, yes.